Find three consecutive even integers such that three times the second equals four more than the sum of the first and third
Answer by richwmiller(17219)

n
n+2
n+4
3*(n+2)=4+n+n+4
3n+6=2n+8
n=2
n+2=4
n+4=6
check
3*4=4+8
12=12
ok